Status of Neutrino Factory and Muon Collider Research and

Development and Future Plans

(Dated: August 5, 2002)

Abstract

We describe the status of the effort to realize a first neutrino factory and the progress made in understanding the problems associated with the collection and cooling of muons towards that end. We summarize the physics that can be done with neutrino factories as well as with intense cold beams of muons. The physics potential of muon colliders is reviewed, both as Higgs Factories and compact high energy lepton colliders. The status and timescale of the research and development effort is reviewed as well as the latest designs in cooling channels including the promise of ring coolers in achieving longitudinal and transverse cooling simultaneously. We detail the efforts being made to mount an international cooling experiment to demonstrate the ionization cooling of muons.

PACS numbers: 13.10.+q, 14.60.Ef, 29.27.-a, 29.20.Dh

[l] Type of  Physics Issues  Possible previously established  present activities  projected for Experiment  Experiments  accuracy  (proposed accuracy)  SMS @ CERN ”Classical”  Lepton Number Violation;  μ^{−}N → e^{−}N  6.1 · 10^{−13}  PSI, proposed BNL (5 · 10^{−17})  < 10^{−18} Rare &  Searches for New Physics:  μ → eγ  1.2 · 10^{−11}  proposed PSI (1 · 10^{−14})  < 10^{−15} Forbidden  SUSY, L-R Symmetry,  μ → eee  1.0 · 10^{−12}  completed 1985 PSI  < 10^{−16} Decays  R-parity violation,.....  μ^{+}e^{−} → μ^{−}e^{+}  8.1 · 10^{−11}  completed 1999 PSI  < 10^{−13} Muon  G_{F}; Searches for New Physics;  τ_{μ}  18 · 10^{−6}  PSI (2x), RAL (1 · 10^{−6})  < 10^{−7} Decays  Michel Parameters  non(V − A)  typ. few 10^{−3}  PSI, TRIUMF (1 · 10^{−3})  < 10^{−4} Standard Model Tests; Muon  New Physics; CPT Tests  g_{μ} − 2  1.3 · 10^{−6}  BNL (3.5 · 10^{−7})  < 10^{−7} Moments  T- resp. CP-Violation  edm_{μ}  3.4 · 10^{−19}e cm  proposed BNL (10^{−24}e cm) < 5 · 10^{−26}e cm in 2nd lepton generation Muonium Fundamental Constants, μ_{μ},m_{μ},α;  M_{HF S}  12 · 10^{−9}  completed 1999 LAMPF  5 · 10^{−9} Spectroscopy Weak Interactions; Muon Charge  M_{1s2s}  1 · 10^{−9}  completed 2000 RAL  < 10^{−11} Muonic Atoms  Nuclear Charge Radii;  μ^{−}atoms  depends  PSI, possible CERN  newnuclear Weak Interactions  (< r_{p} >to 10^{−3})  structure Condensed  surfaces, catalysis  surface μSR  n/a  PSI, RAL (n/a)  highrate Matter  bio sciences ... TABLE I: Experiments which could beneficially take advantage of the intense future stopped muon source. The numbers were worked out for scenarios at a future Stopped Muon Source (SMS) of a neutrino factory at CERN [? ]. They are based on a muon flux of 10^{21} particles per annum in which beam will be available for 10^{7} s. Typical beam requirements are given in Table ??.

2